The graph of the fns y=f(x) is symmetrical about the line x=2 then



a f(x+2)=f(x-2)

b f(2+x)=f(2-x)

c f(x)=f(-x)

d f(x)=-[f(-x)]

Grade:12

3 Answers

Sunil Kumar
10 years ago

Answer : (b) f(2+x)=f(2-x)

 check for thr conditions of symmetry.

 

Thanks & Regards
Sunil Kumar, 
askIITians Faculty
M.Sc., IIT Kharagpur

Himanshu Mohan
10 years ago
can u give me solution plz
ayush kumar
10 years ago
since function is symmetric about x=2; value of function will be identical on both side of 2 or we can say x=2 is a mirror . so whatever distance you move from x=2 in right means you are moving same distance in left. i.e. if you are moving x in the right you are also moving x in the left and therefore these two will have same value. i.e. f(2+x)=f(2-x)